Q:

A couple wishes to replace the carpeting in their master bedroom. Their master bedroom measures 24 ft by 16 ft. If carpeting is sold by the square yard, how many square yards of carpeting are needed?


1
Expert's answer
2022-01-14T06:51:36-0500

1yd=3ft, so 1ft=0.33yd

1ft=0.33yd

24ft=?yd

(24ft/1ft)*0.33yd= 8yd

16ft=?yd

16ft/1ft*0.33yd= 5.33yd

"Area= L*W"

"=8*5.33" "=42.64yd^2"


alternatively

"24ft*16ft=384ft^2"

"9ft^2=1yd^2"

"(384ft^2\/9ft^2)*1yd^2"

"=42.64yd^2"

square yards of carpeting that are needed is "42.64yd^2"